OpenStudy (anonymous):

Solve the equation 3.7(-83x-3)=3.7(7x-30)

jigglypuff314 (jigglypuff314):

to solve for x: 3.7(-83x-3)=3.7(7x-30) divide by 3.7 on both sides /3,7 /3.7 (-83x-3) = (7x-30) then add 30 to both sides +30 +30 -83x + 27 = 7x + 0 then add 83x to both sides +83x +83x 0 + 27 = 90x then divide both sides by 90 /90 /90 27/90 = x then you can simplify x = 3/10